Solve. x^2 + 3 = 39
I came up with +-SQRT26
Am I right?

Answer by dolly(163)


No, You aren't.
x^2 + 3 = 39
==> x^2 = 39 - 3
==> x^2 = 36
==> x = +/- 6
==> x = 6, -6

Good Luck!!!